OpenStudy (av0cad0):

x/3 + x/2 = 5 I don't know how to get started.

OpenStudy (adamaero):

Do you like fractions?

OpenStudy (av0cad0):

I really don't actually. I've never been good with them...

OpenStudy (adamaero):

Then get rid of them! You can either do this in one blow or you can do it for each.

OpenStudy (adamaero):

Multiply everything by 3. Then multiply everything by 2. (Or just multiply everything by 2*3)

OpenStudy (3mar):

I like your expressions, @adamaero Thanks for good illustration!

OpenStudy (hoot_hoot):

Find the LCD of both fractions. Divide the LCD by each denominator then multiply it to the numerator.

OpenStudy (mathmale):

The purpose of all these goings-on with LCDs is to clear the given equation of fractions. What is the LCD? Multiply all three terms of the given equation by this LCD. Reduce each term as necessary. The resulting equation will have no fractions in it. Show your work, please.

OpenStudy (mathmale):

Would you please stick around for at least a few minutes, in case someone begins to help you?

OpenStudy (naka354):

x*(1/3 + 1/2) = 5 x*(5/6) = 5 x = 5/(5/6) x = 6